Find a general solution of (4D^2+4D+17)y= 0

1
Expert's answer
2021-12-15T12:24:58-0500

The auxiliary equation is

"4m^2+4m+17=0"

"m = \dfrac{-4 \pm \sqrt{16-68}}{8}"

"m = \dfrac{-2 \pm \sqrt{13}i}{4}"

The general solution of the differential equation is

"y = e^{-\frac{1}{2}t}[A\cos( \frac{\sqrt{13}}{4}t) + iB \sin( \frac{\sqrt{13}}{4}t)]"